OpenStudy (anonymous):

Solve for x. 4 • 7 – 4 • 10 = 4(x – 10) A. –13 B. –6 C. 4 D. 7

OpenStudy (jadeishere):

4*7 - 4 * 10 = 4x - 40 28 - 40 = 4x - 40 +40 + 28 - 40 = 4x 28 = 4x x = 7

OpenStudy (anonymous):

28-40=4(x-10)